# Learning normalized image densities via dual score matching

### v1

@@ 2.3 Dual score matching: Architecture
How should one choose an architecture to compute the energy $U_{\theta}(y,t)$? Rather than designing one from scratch (Salimans and Ho, 2021; Cohen et al., 2021; Thiry and Guth, 2024), we instead construct one by modifying a score-based denoising architecture that is known to have appropriate inductive biases. Let $s_{\theta}\colon\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ be such an architecture (e.g., a UNet). We wish to define a new architecture $U_{\theta}\colon\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}$ such that $\nabla_{y}U_{\theta}\approx s_{\theta}$, preserving the inductive biases of $s_{\theta}$. We propose to set
@@
We show in Section B.2 that $\nabla_{y}U_{\theta}(y,t)=s_{\theta}(y,t)$ if the score network $s_{\theta}$ is conservative and homogeneous (Romano et al., 2017; Reehorst and Schniter, 2018). Homogeneity has been shown to hold approximately in the related setting of blind denoisers (Mohan* et al., 2020; Herbreteau et al., 2023). We note that the seemingly similar choice of a squared norm such as $U_{\theta}(y,t)=\frac{1}{2}{\mathopen{}\mathclose{{}\left\lVert s_{\theta}(y,t)}\right\rVert}^{2}$ used in some previous energy models (Salimans and Ho, 2021; Hurault et al., 2021; Du et al., 2023; Yadin et al., 2024; Thiry and Guth, 2024) leads to the same desirable homogeneity properties in $y$ but not in $\theta$, and thus fails to preserve the optimization properties of the original score network. We describe further architectural details in Section C.1.
@@ 2.4 Performance evaluation
We first verify that the gradient of the energy network $U_{\theta}$ provides as good a denoiser as the score network on which it is based, $s_{\theta}$. Both networks are separately trained; $U_{\theta}$ with the dual score matching objective (eq. 8, using double back-propagation) and $s_{\theta}$ with only the standard (denoising) score matching objective (eq. 3). Further details are provided in Section C.2. We compare their denoising performance across noise levels in Table 1. For all but the smallest noise levels, the energy-based model achieves (slightly) better denoising performance than the score model. These results show that there is no penalty in modeling the energy rather than the score. This is contrary to the results of Salimans and Ho (2021), and is likely due to our use of an architecture that is homogeneous, and for which non-conservativeness of the score is not a large source of denoising error (Mohan* et al., 2020; Chao et al., 2023). Finally, these results also demonstrate that the two components of our dual score matching objective are not trading off against each other; rather, they complement and even reinforce each other.
Table 1: Denoising MSE at several noise levels for corresponding score and energy networks, averaged over images in ImageNet64. All quantities are expressed as peak signal-to-noise ratio (PSNR) in dB: $\mathrm{PSNR}=-10\log_{10}(\mathrm{MSE})$.
Noise variance $90$ $75$ $60$ $45$ $30$ $15$ 00 $-15$ $-30$
Score network $\mathbf{90.20}$ $\mathbf{75.47}$ $60.43$ $47.19$ $35.58$ $25.92$ $18.84$ $13.31$ $-0.11$
@@
How accurate are our energy estimates? A standard evaluation of probabilistic models consists of calculating the cross-entropy $\mathbb{E}_{x}\mathopen{}\mathclose{{}\left[-\log p_{\theta}(x)}\right]$ between the data distribution and the model, also known as negative log likelihood (NLL). In practice, NLL computes how much probability the model assigns to a held-out test set. There are however several subtleties that make direct comparisons challenging. Specifically, there are 3 factors that can cause variations in NLL up to ±$\pm 0.5$ bits/dimension or more: details of data pre-processing (e.g., downsampling or data augmentation), conversion method from continuous to discrete probabilities, and the estimator type (exactly normalized, variational bound, or approximately normalized). We expand on these issues in Section A.1.
We compare NLLs of our method and a variety of recent energy models in Table 2. This evaluation demonstrates that our model is comparable to the best-performing models in the literature, within the variability arising from the three factors of variation mentioned in the previous paragraph. Two unique advantages of our method are that it provides (1) direct (one-shot) estimates of energy (as opposed to other density estimation approaches in diffusion models, which we review in Section A.2), greatly facilitating analysis, and (2) access to energy across all noise levels, providing a window into larger-scale features of the energy landscape.
Table 2: Negative log likelihood (in bits/dimension) on ImageNet64 test set. See Section A.1 for more details.
@@ 3.1 Generalization
Figure 2: Convergence of energy estimates. The data set is split into two halves (denoted A and B), and separate energy models are trained on $N$ samples drawn from each half. Each scatterplot compares the $\log p(x)$ estimated by the two models at $t=0$ over all $2N$ training images. As $N$ increases, the energy estimates of the two models converges for all images.
The previous section demonstrated that our energy-based model achieves near state-of-the-art NLL on ImageNet64. That is, the model on average assigns high probability to a set of held-out images. Next, we establish that the energies of the individual images are reliable. In particular, we require that the model’s energy assignment is stable under change of the training data.
To this end, we borrow the strong generalization test developed in Kadkhodaie et al. (2024). We partition the training data into two non-overlapping sets, train a separate energy-based model on each set, and then compare the energies computed by these two models on their own and each other’s training examples. We gradually increase the size of each training set until the two models assign approximately equal log probability across all images. Figure 2 shows the results of this experiment. The two models assign very different probabilities to the same image when the training set size, $N$, is small. But they converge gradually and compute almost the same values at $N=10^{5}$. Note that the rate of convergence depends on the probability: more data is needed before the two models agree on the high-probability images. It is also worth noting that the transition from memorization to generalization of energy models is marked by a large increase of variance in energy over the training set (starting at $N=100$).
This result establishes that the model variance vanishes with a feasible training set size, demonstrating lack of memorization. However, it does not guarantee that the values are accurate, as the models could be biased. Calculating model bias is not possible without access to the true density, hence it can only be obtained for simple analytical models (as in Figure 1). However, smaller NLL values over test data (Table 2) are indicative of a smaller model bias.
@@ 3.2 Distribution of energies and relationship to image content
Figure 3: Histogram of log probabilities of images in the ImageNet dataset. Color-coded arrows indicate values for the example images on the right, and the leftmost (brown) and rightmost (green) arrows indicate values for a uniform noise image in $[0,1]$ and a constant image of intensity $0.5$, respectively. The distribution is well-fit by a Gumbel distribution (red line). Additional examples of images organized by probability are shown in Figures 6, 7 and 8 (Appendix D).
We now study properties of the learned energy model. What is the entropy (average energy) of the image dataset? Do all images have the same probability? If not, what determines the probability of an image? To investigate these questions, we compute the log probability of all $50,000$ images in the ImageNet test set in Figure 3. We express log probability in units of decibels per dimension (dB/dim), computed as $\frac{1}{d}10\log_{10}p_{\theta}(x)$. In this scale, an additive change of $10$ dB/dim corresponds to a multiplicative change in probability density of $10^{d}$.
@@
The average of $-\log p_{\theta}(x)$ provides an estimate of the differential entropy of ImageNet, which is here equal to $-11.4$ dB/dim. Note that the uniform distribution over the hypercube $[0,1]^{d}$ has an entropy of 00. This indicates that natural images occupy a fractional volume of about $10^{-1.14d}$. This can be converted to an estimate of discrete entropy by assuming that log probability is constant within quantization bins. For $256$ possible intensity values ($8$ bits), this gives an entropy of $4.20$ bits/dimension (the difference with Table 2 comes from the use of grayscale images here). In other words, there are $\sim 10^{5,180}$ quantized ImageNet images out of $10^{9,860}$ possible images at this resolution.
@@
Many high-dimensional probability distributions exhibit a concentration phenomenon, where typical realizations have nearly equal energy. For instance, in simple image probability models such as a Gaussian model or a sparse wavelet model (where wavelet coefficients are independent), the energy of independent components add, and the law of large numbers imply a concentration of the energy around its mean. In contrast, our energy network reveals enormous diversity in the log probabilities of individual ImageNet images (Figure 3). Over the entire test set, they span a range of $34.4$ dB/dim, corresponding to a probability ratio of $\approx 10^{14,000}$. To make sense of this value, it is important to realize that the frequency of images with a given value of $\log p$ in a dataset is equal to this probability value multiplied by the volume of the corresponding probability level set. Observing a range of probability values of $10^{14,000}$ thus reveals that these enormous variations in probability density are nearly compensated by inverse variations in the volume of their corresponding level sets.
@@
The distribution of log probabilities values is highly skewed: there is a heavy tail of high-probability images and a much lighter tail of low-probability images. Surprisingly, this distribution is well approximated by a Gumbel distribution (we report parameter values in Section C.3). It arises as a limit distribution for the maximum of many i.i.d. random variables. This surprising observation is, to the best of our knowledge, new, and calls for an explanation. In independent component models, the log probabilities are sums of i.i.d. random variables, and are therefore Gaussian distributed (with a vanishing variance compared to their mean). A simple model which reproduces this Gumbel distribution is a high-dimensional spherically-symmetric distribution with an exponentially-distributed radial marginal.
Figure 4: Influence of image statistics on probability. Left. $\log p_{\theta}(ax+b)$ as a function of $a$ and $b$. Middle. Horizontal slice ($b=\frac{1}{2}$) of the left plot for the example images of Figure 3. Right. Log probability as a function of sparsity, measured as the participation ratio of wavelet coefficients.
@@
We also show in Figure 3 examples of images at various log probabilities. High-probability images invariably contain small (often high-contrast) objects, on a blank (often white) background. Conversely, low-probability images are generally filled with dense detailed texture. This agrees with the behavior of compression engines. Indeed, the energy of an image, when expressed in bits, corresponds to the size of its optimally compressed representation. This implies that low-probability images should intuitively have more “content” than high-probability images.
@@
Following these visual observations, we further examine the influence of intensity range and sparsity on image probability (Figure 4). First, we evaluate the log probability of reference images as we manipulate their brightness or contrast through an affine operation. We find that the brightness has minimal effect on the probability, while higher-contrast images have lower probability. This reveals that the distribution is star-shaped: any pair of images are connected by a high-probability path passing through a constant image. Next, we verify that log probability is correlated with a simple measure of sparsity of images. We compute a multi-scale wavelet decomposition of the images, and measure the $\ell^{1}$-norm divided by the $\ell^{2}$-norm of the coefficients. The square of this quantity, $\|x\|_{1}^{2}/d\|x\|^{2}_{2}\in(0,1]$ is known as participation ratio, with smaller values indicating higher sparsity. The right panel of Figure 4 shows that this simple measure is indeed predictive of $\log p_{\theta}(x)$ (albeit not very precisely).
@@ 3.3 Effective dimensionality of the energy landscape
Beyond estimating log probability for a given image, we aim to characterize the local behavior of the density in the vicinity of that image. For instance, we would like to assess whether the probability is locally concentrated near a low-dimensional “tangent” subspace, and if so, estimate its dimensionality. In this section, we explain how these quantities may be computed from our energy model. We then show that the local dimensionality around an image is often much lower than the ambient dimensionality of the space, but that this dimensionality varies with the particular image and the scale that is used to define the local neighborhood.
@@
Consider the distribution supported on the blue regions in the left two panels of Figure 5. For the leftmost region, the effective dimensionality of a local neighborhood decreases with the size of that neighborhood. The opposite behavior is also possible, as shown in the rightmost region. These examples show that dimensionality measures which aim to describe these geometrical structures need to depend on both the location and the scale of the neighborhood (Tempczyk et al., 2022).
@@
We show in Section B.3 that the definitions in eqs. 11 and 12 are equivalent, unifying two seemingly distinct points of views adopted in related dimensionality measures (Wu and Verdú, 2011; Mohan* et al., 2020; Tempczyk et al., 2022; Stanczuk et al., 2022; Kamkari et al., 2024; Horvat and Pfister, 2024).
@@
Figure 5: Left: Two hypothetical examples illustrating how local effective dimensionality depends on the scale of the neighborhood. For both examples, support of density corresponds to the blue regions. In the left example, the dimensionality around the red point decreases with scale (from 2 down to 0), while the opposite is true for the right example. Right: Log probability and effective dimensionality as a function of noise level. Colored lines correspond to different example images $x$ (shown in right panel of Figure 3), while the dashed black line shows the average over the ImageNet test set. The vertical gray line indicates the minimum noise level presented during training ($t=10^{-9})$, and the horizontal gray line the ambient dimensionality of the dataset ($d=4096$).
@@
We show in Figure 5 the behavior of the effective energy and dimensionality as a function of the noise level. As the noise level increases, the log probability initially remains constant, indicating that probability is uniformly spread in a ball around $x$. The probability eventually decreases as the diffusion radius becomes larger than the local support size, and all lines converge to the mean (the negative entropy), consistent with the asymptotic Gaussian behavior of the energy. The effective dimensionality vanishes at large noise levels (images have a compact support, which eventually looks like a point at large enough scales) but increases as the noise level is reduced, eventually reaching the ambient dimensionality $d$ (our model has a continuous non-zero density everywhere by construction). In-between these two extremes, there exists a sizable range of scales $t\in[{10}^{-9},{10}^{-2}]$ where almost the entire range of dimensionalities coexist. In particular, higher-probability images have lower-dimensional neighborhoods, even at relatively small scales, while lower-probability images have nearly full-dimensional neighborhoods, even at relatively large scales. We note two potential limits to these estimates. First, the quantization of pixel values into bins of size $1/256$ limits the resolution to $t\sim{10}^{-5}$, possibly explaining the momentary decrease in dimensionality as $t$ decreases. Second, our energy model was only trained on noise levels down to $t_{\min}={10}^{-9}$. Thus, energies at smaller values of $t$ are solely determined by model inductive biases, which favor a constant energy.

### v3

@@ 2.3 Dual score matching: Architecture
How should one choose an architecture to compute the energy $U_{\theta}(y,t)$? Rather than designing one from scratch (Salimans and Ho, 2021; Cohen et al., 2021; Thiry and Guth, 2024), we construct one by modifying a score-based denoising architecture that is known to have appropriate inductive biases. Let $s_{\theta}\colon\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ be such an architecture (e.g., a UNet). We wish to define a new energy architecture $U_{\theta}\colon\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}$ such that $\nabla_{y}U_{\theta}\approx s_{\theta}$, preserving the inductive biases of $s_{\theta}$. To achieve this, we set
@@
This inner product parameterization has been concurrently proposed by Thornton et al. (2025). We show in Section˜B.2 that $\nabla_{y}U_{\theta}(y,t)=s_{\theta}(y,t)$ if the score network $s_{\theta}$ is both conservative and homogeneous (Romano et al., 2017; Reehorst and Schniter, 2018). Homogeneity has been shown to hold approximately in the related setting of blind denoisers (Mohan* et al., 2020; Herbreteau et al., 2023), and can be enforced architecturally. Note that the seemingly similar choice of a squared norm architecture $U_{\theta}(y,t)=\frac{1}{2}{\mathopen{}\mathclose{{\left\lVert s_{\theta}(y,t)}}\right\rVert}^{2}$ used in some previous energy models (Salimans and Ho, 2021; Hurault et al., 2021; Du et al., 2023; Yadin et al., 2024; Thiry and Guth, 2024) leads to the same desirable homogeneity properties in $y$ but not in $s_{\theta}$, and thus fails to preserve the optimization properties of the original score network. Further architectural details are provided in Section˜C.1.
@@ 2.4 Performance evaluation
We first verify that the gradient of the proposed energy network, $\nabla_{y}U_{\theta}$, provides as good a denoiser as the score network $s_{\theta}$ on which it is based. We train the two networks separately, $U_{\theta}$ with the dual score matching objective (eq.˜8 using double back-propagation), and $s_{\theta}$ with the standard (denoising) score matching objective (eq.˜3). Further details are provided in Section˜C.2. Table˜1 compares denoising performance across noise levels. For all but the smallest noise levels, the energy-based model achieves (slightly) better denoising performance than the score model. Thus, there is no penalty in modeling the energy rather than the score. This is contrary to the results of Salimans and Ho (2021), and is likely due to our use of an architecture that is homogeneous, and for which non-conservativeness of the score is not a large source of denoising error (Mohan* et al., 2020; Chao et al., 2023). Finally, these results demonstrate that the two components of our dual score matching objective are not trading off against each other; rather, they complement and even reinforce each other (i.e. their minima coincide).
Table 1: Denoising performance at several noise levels for corresponding score and energy networks, averaged over images in ImageNet64, with intensities scaled to lie in $[0,1]$. All quantities are expressed as peak signal-to-noise ratio (PSNR) in dB: $\mathrm{PSNR}=-10\log_{10}(\mathrm{MSE})$.
Noise variance $90$ $75$ $60$ $45$ $30$ $15$ 0 15-15 30-30
Score network $\mathbf{90.20}$ $\mathbf{75.47}$ $60.43$ $47.19$ $35.58$ $25.92$ $18.84$ $13.31$ 0.11-0.11
@@
How accurate are our energy estimates? A standard evaluation of probabilistic models consists of estimating the cross-entropy $\mathbb{E}_{x}\mathopen{}\mathclose{{\left[-\log p_{\theta}(x)}}\right]$ between the data distribution and the model, also known as negative log likelihood (NLL). In practice, NLL is computed by computing the probability the model assigns to a held-out test set. There are however several subtleties that make direct model comparisons challenging. Specifically, there are 3 factors that can cause variations in NLL of $\pm 0.5$ bits/dimension or more: details of data pre-processing (e.g., downsampling or data augmentation), conversion method from continuous to discrete probability, and the estimator type (exactly normalized, variational bound, or approximately normalized). We expand on these issues in Section˜A.1.
We compare NLLs of our method and a variety of recent energy models in Table˜2. This evaluation demonstrates that our model is comparable to the best-performing models in the literature, within the variability arising from the three factors of variation mentioned in the previous paragraph. Two unique advantages of our method are that it provides (1) direct (one-shot) estimates of energy (as opposed to other density estimation approaches used with diffusion models, which we review in Section˜A.2), and (2) access to energy across all noise levels, providing a window into larger-scale features of the energy landscape. For instance, our energy network can compute the probability of 50k images in 12s on an A100 GPU, whereas the score network requires upwards of 3h20, even with as few as 100 times steps and 10 noise samples to compute the energy (see Section˜A.2). This justifies the longer training time due to the cost of double-backpropagation (training for 1M steps on ImageNet64 on a single A100 GPU respectively took 120 hours, compared to 32 hours for the score network), which may be alleviated by the use of sliced score matching (Song et al., 2020).
Table 2: Negative log likelihood (in bits/dimension) on ImageNet64 test set. See Section˜A.1 for more details.
@@ 3.1 Generalization
Figure 2: Convergence of energy estimates. The data set is split into two halves (denoted A and B), and separate energy models are trained on NN samples drawn from each half. Each scatterplot compares the energy estimates of the two models at $t=0$, over all $2N$ training images. As NN increases, the energy estimates of the two models converges for all images.
The previous section demonstrated that our energy-based model achieves near state-of-the-art NLL on ImageNet64. That is, the model on average assigns high probability to a set of held-out test images. Next, we establish that the energies of the individual images are reliable. In particular, we verify that the model’s energy assignment is stable under change of the training data.
To this end, we use the strong generalization test developed in Kadkhodaie et al. (2024). We partition the training data into two non-overlapping sets, train a separate energy-based model on each set, and then compare the energies computed by these two models on images from both training subsets. We gradually increase the size of each training set until the two models assign approximately equal log probability across all images. Figure˜2 shows the results of this experiment. The two models assign very different probabilities to the same image when the training set size, $N$, is small. But they converge gradually and compute nearly the same values at $N=10^{5}$. Note that the rate of convergence depends on image probability: more data is needed before the two models agree on the high-probability images. It is also worth noting that the transition from memorization to generalization of energy models is marked by a large increase of variance in energy over the training set (starting at $N=100$).
This result establishes that the model variance vanishes with a feasible training set size. However, it does not guarantee that the values are accurate—the models could be biased. Direct calculation of model bias requires access to the true density, which is only available for synthetic data (as in Figure˜1). However, the small NLL values over test data (Table˜2) do suggest a small model bias.
@@ 3.2 Distribution of energies and relationship to image content
Figure 3: Histogram of log probabilities of test images in the ImageNet dataset. Color-coded arrows indicate values for the example images on the right, and the leftmost (brown) and rightmost (green) arrows indicate values for a uniform noise image in $[0,1]$ and a constant image of intensity $0.5$, respectively. The distribution is well-fit by a Gumbel distribution (red line). Additional example images organized by probability are provided in Figures˜6, 7 and 8 (Section˜D.1).
We now study properties of the learned energy model. What is the entropy (average energy) of the image dataset? Do all images have the same probability? If not, what determines the probability of an image? To investigate these questions, we compute the log probability of all $50,000$ images in the ImageNet64 test set. We express log probability in units of decibels per dimension (dB/dim), computed as $\frac{1}{d}10\log_{10}p_{\theta}(x)$, with $d=64\times 64=4096$. In this scale, an increase of $10$ dB/dim corresponds to a multiplicative rescaling of the probability density by a factor of $10^{d}$.
@@
The average value of $\log p_{\theta}(x)$ provides an estimate of 11.4-11.4 dB/dim for the differential entropy of ImageNet. Note that the uniform distribution over the hypercube $0,1]^{d}$ has an entropy of 0. This indicates that natural images occupy a fractional volume of about $10^{-1.14d}$. This can be converted to an estimate of discrete entropy by assuming that log probability is constant within quantization bins. For $8$-bit images (with $256$ possible intensity values), this corresponds to an entropy of $4.20$ bits/dimension (the deviation from the value in Table˜2 is due to the use of grayscale images here). In other words, there are $\sim 10^{5,180}$ quantized ImageNet images out of $10^{9,860}$ possible images at this resolution.
@@
Many high-dimensional probability distributions exhibit a concentration phenomenon: typical realizations have nearly equal energy. For instance, in simple image probability models such as a Gaussian model or a sparse wavelet model (where wavelet coefficients are independent), the energy of independent components add, and the law of large numbers imply a concentration of the energy around its mean. In contrast, our energy network reveals enormous diversity in the log probabilities of individual ImageNet images (Figure˜3, left). Over the entire test set, they span a range of $34.4$ dB/dim, corresponding to a probability ratio of $\approx 10^{14,000}$. Given this enormous ratio, it is surprising that any low probability image appears in a finite dataset. To make sense of this apparent contradiction, it is important to realize that the total probability mass of images with a given value of $\log p$ in a dataset is equal to this probability value multiplied by the volume of the corresponding probability level set. Observing a range of probability values of $10^{14,000}$ thus reveals that these enormous variations in probability density must be nearly compensated for by inverse variations in the volume of their corresponding level sets.
@@
The distribution of log probability values is highly skewed (Figure˜3, left): there is a heavy tail of high-probability images and a much lighter tail of low-probability images. We also observed these qualitative features on the CelebA and CIFAR-10 datasets, shown in Section˜D.2. Surprisingly, this distribution is well approximated by a Gumbel distribution (we report parameter values in Section˜C.3), which arises as a limit distribution for the maximum of many i.i.d. random variables. This surprising observation is, to the best of our knowledge, novel, and calls for an explanation. In independent component models, the log probabilities are sums of i.i.d. random variables, and are therefore Gaussian-distributed in high dimensions. In contrast, a simple model that leads to a Gumbel distribution is a high-dimensional spherically-symmetric distribution with an exponentially-distributed radial marginal (Lyu and Simoncelli, 2009).
Figure 4: Influence of image statistics on probability. Left. $\log p_{\theta}(ax+b)$ as a function of $a$ and $b$. Middle. Horizontal slice ($b=\frac{1}{2}$) of the left panel for the example images of Figure˜3. Right. Log probability as a function of sparsity, measured as the participation ratio of wavelet coefficients.
@@
We also show in Figure˜3 (right) example images with a variety of log probabilities. High-probability images contain small objects, on a blank (often white) background. Conversely, low-probability images are filled with dense detailed texture. Since the energy of an image, when expressed in bits, corresponds to the size of its optimally compressed representation, this implies that higher-probability images have shorter codes and thus less “information content”. We show in Section˜D.4 images that have been obtained by locally maximizing or minimizing probability.
@@
Following these visual observations, we further examine the influence of intensity range and sparsity on image probability (Figure˜4). First, we evaluate the log probability of reference images as we manipulate their brightness and contrast through an affine operation. We find that the brightness has minimal effect on the probability, while higher-contrast images have lower probability. This reveals that the distribution support is star-shaped: any pair of images are connected by a high-probability path passing through a constant image. In contrast, we show in Section˜D.3 that linear interpolation between pairs of images typically increase probability: the distribution support is not convex. Next, we verify that log probability is correlated with a simple measure of sparsity of images. We compute a multi-scale wavelet decomposition of the images, and measure the $\ell^{1}$-norm divided by the $\ell^{2}$-norm of the coefficients. The square of this quantity, $\|x\|_{1}^{2}/d\|x\|^{2}_{2}\in(0,1]$ is known as the participation ratio, with smaller values indicating higher sparsity. The right panel of Figure˜4 shows that this simple measure captures a significant portion of the variance of $\log p_{\theta}(x)$.
@@ 3.3 Effective dimensionality of the energy landscape
Beyond estimating log probability for a given image, we aim to characterize the local behavior of the density in the vicinity of that image. For instance, we would like to assess whether the probability is locally concentrated near a low-dimensional “tangent” subspace, and if so, to estimate its dimensionality. In this section, we explain how these quantities may be computed from our energy model. We then show that the local dimensionality around an image is often much lower than the ambient dimensionality of the space, but that this dimensionality depends on both the particular image and the scale that is used to define the local neighborhood.
@@
Consider hypothetical distributions supported on the blue regions in the left two panels of Figure˜5. For the left example, the effective dimensionality of a local neighborhood decreases with the size of that neighborhood. The opposite behavior is also possible, as shown in the right example. These examples demonstrate that dimensionality measures which aim to describe these geometrical structures need to depend on both the location and the scale of the neighborhood (Mohan* et al., 2020; Tempczyk et al., 2022).
@@
We show in Section˜B.3 that the definitions in eqs.˜11 and 12 are equivalent, unifying two seemingly distinct points of view underlying these dimensionality measures (Wu and Verdú, 2011; Mohan* et al., 2020; Tempczyk et al., 2022; Stanczuk et al., 2022; Kamkari et al., 2024; Horvat and Pfister, 2024).
@@
Figure 5: Left: Two hypothetical examples illustrating how local effective dimensionality depends on the scale of the neighborhood. For both examples, support of density corresponds to the blue regions. In the left example, the dimensionality around the red point decreases with scale (from 2 down to 0), while the opposite is true for the right example. Right: Log probability and effective dimensionality as a function of noise level. Colored lines correspond to different example images $x$ (shown in right panel of Figure˜3), while the dashed black line shows the average over the ImageNet test set. The vertical gray line indicates the minimum noise level presented during training ($t=10^{-9})$, and the horizontal gray line the ambient dimensionality of the dataset ($d=4096$).
@@
We show in Figure˜5 the behavior of the effective energy and dimensionality as a function of the noise level. As the noise level increases, the log probability initially remains constant, indicating that probability is uniformly spread in a ball around $x$. The probability eventually decreases as the diffusion radius becomes larger than the local support size, and all lines converge to the mean (the negative entropy), consistent with the asymptotic Gaussian behavior of the energy. The effective dimensionality (fig.˜5, right) vanishes at large noise levels (images have a compact support, which behaves like a point at sufficiently large noise levels) but increases as the noise level is reduced, eventually reaching the ambient dimensionality $d$ (our model has a continuous non-zero density everywhere by construction). In-between these two extremes, there exists a sizable range of scales $t\in[{10}^{-9},{10}^{-2}]$ where almost the entire range of dimensionalities coexist. In particular, higher-probability images have lower-dimensional neighborhoods, even at relatively small scales, while lower-probability images have nearly full-dimensional neighborhoods, even at relatively large scales. This result empirically confirms the point raised in Section˜3.2: the local probability mass around high and low probability images are of the same order, despite the enormous gap between their probability values. The local high dimensionality of the low-probability images corresponds to a significantly larger volume occupied by their local neighborhood. We note two potential limits to these estimates. First, the quantization of pixel values into bins of size $1/256$ limits the resolution to $t\sim{10}^{-5}$, possibly explaining the momentary decrease in dimensionality as $t$ decreases. Second, our energy model was only trained on noise levels down to $t_{\min}={10}^{-9}$. Thus, energies at smaller values of $t$ are solely determined by model inductive biases, which favor a constant energy.